OpenStudy (anonymous):

f(x)=(1x-9)/(14x+7); f^-1(x)=?

OpenStudy (anonymous):

ho ho ho a raft of algebra ready ?

OpenStudy (anonymous):

\[x=\frac{y-9}{14y+7}\] solve for \(y\) first clear the fraction \[x(14y+7)=y-9\] multiply out \[14xy+7x=y-9\] all \(y\) on the left everything else on the right \[14xy-y=-9-7x\] factor \[(14x-1)y=-9-7x\] divide \[y=\frac{-9-7x}{14x-1}\]

OpenStudy (anonymous):

Thanks, that helped a lot!

OpenStudy (anonymous):

yw